A ball of mass 625.g traveling 4.00 m/s to the right, has an elastic head-on collision with an identical ball which is at rest. Find the final velocity of each ball.

1
Expert's answer
2022-02-19T04:55:54-0500

Given:

"m_1=m_2=625\:\rm g"

"v_1=4.00\;\rm m/s"

"v_2=0\;\rm m/s"


The final velocities of balls

"v_1'=\frac{m_1-m_2}{m_1+m_2}v_1=0\:\rm m/s"

"v_2'=\frac{2m_1}{m_1+m_2}v_1=v_1=4.00\:\rm m/s"